OpenStudy (anonymous):

Divide and check (x2 + 2x − 15) ÷ (x + 5)

OpenStudy (callisto):

(x^2 + 2x − 15) ÷ (x + 5) = (x+5)(x-3) ÷ (x + 5) \[=\frac{(x+5)(x-3)}{(x+5)}\] Can you cancel the common factor now?

OpenStudy (anonymous):

Yes I got it , thanks

OpenStudy (callisto):

Welcome!~